OpenStudy (anonymous):

A spring stretches linearly as weight is added. The table shows data collected for a certain spring. What is the slope of the line that fits the data for the spring? Weight (g) .... 100 500 800 900 1,200 Stretch (cm) ... 0.5 2.5 4.0 4.5 6.0 a. 1/200 b. 1/100 c. 1/50 d. 1/2

OpenStudy (anonymous):

y=kx y be the wieght and x stretch sry weight k = 5/10000 k= 1/200